What is the speed of a car that travels 250m in 55s? If you can explain rather than just give the answer that would be great. I have a final tomorrow.
Answer by Alan3354(69443)

What is the speed of a car that travels 250m in 55s?
----------------
d = r*t
r = d/t
r = 250m/55s
r = 50/11 m/sec
r =~ 4.545 m/sec
==================
If you want other units, you'll have to specify.